OpenStudy (zyberg):

a, b, c, d - positive numbers. Prove that there is a triangle that has sides of such lengths: square root of (b^2 + c^2), square root of (a^2 + c^2 + d^2 + 2cd), square root of (a^2 + b^2 + d^2 + 2ab) and find the area of this triangle.

OpenStudy (zyberg):

\[\sqrt{b^2 + c^2}, \sqrt{a^2 + c^2 + d^2 + 2cd}, \sqrt{a^2 + b^2 + d^2 + 2ab}\]

OpenStudy (zyberg):

I notice that the second length can be simplified to: \(\sqrt{a^2 + (c +d)^2}\) and the third length to: \(\sqrt{d^2 + (a + b)^2}\)

OpenStudy (zyberg):

However, I have no idea what to do with this knowledge.

OpenStudy (samigupta8):

What's the essential thing for a triangle to be made?

OpenStudy (zyberg):

Perhaps a base? Or the height? (Not sure how it is called in English, can't find the translation, the thing that is usually denoted with a letter h)

OpenStudy (samigupta8):

Hypotenuse!

OpenStudy (zyberg):

But what do I get from that knowledge?

OpenStudy (samigupta8):

Lol! No , i am asking for a particular property that guides the formation of a triangle.

OpenStudy (samigupta8):

You must be familiar with triangle inequality.

OpenStudy (zyberg):

Well, I think that you are talking about the idea that in atriangle there are two or one triangles with one angle of 90 degrees? So it is possible to apply Pythagorean theorem? I am familiar with it, but how would it help there?

OpenStudy (samigupta8):

Okay! Let's start with basics .

OpenStudy (samigupta8):

http://mathworld.wolfram.com/TriangleInequality.html

OpenStudy (samigupta8):

Have a look at it.

OpenStudy (zyberg):

So, first I would take any of the lengths, make up such inequality, square both sides and look what goes there? But since I don't know the numeric values of letters, how would I know, if the inequality holds? (I need to prove that those lengths really make up a triangle)

OpenStudy (samigupta8):

We don't need that rather . There will be a lot much simplification after you do that process .

OpenStudy (zyberg):

I will try to do that: \(\sqrt{a^2 + b^2} < \sqrt{(c+d)^2 + a^2 }+\sqrt{(a+b)^2 + d^2}\) Squaring both sides: \(a^2 + b^2 < (c +d)^2 + a^2 + 2(\sqrt{(c+d)^2 + a^2 })(\sqrt{(a+b)^2 + d^2}) + (a +b)^2 +d^2\)

OpenStudy (zyberg):

So: \(0 < (c+d)^2 + 2(\sqrt{(c+d)^2 + a^2 })(\sqrt{(a+b)^2 + d^2}) + a^2 + 2ab +d^2\)

OpenStudy (samigupta8):

I think that you did a mistake .

OpenStudy (zyberg):

Where?

OpenStudy (samigupta8):

It was square root of b^2+c^2 and not of a^2+b^2.

OpenStudy (zyberg):

Ahh! My bad. Got ahead of myself a little bit with those letters ;)

OpenStudy (samigupta8):

Happens!

OpenStudy (zyberg):

But exchanging last terms of RHS a to c, and we got proof? Or do I need to do that for all sides?

OpenStudy (samigupta8):

The first half would do ..

OpenStudy (samigupta8):

Sorry! It's getting l8 here.. I am really sleepy at the moment.

OpenStudy (zyberg):

No problem! You can go sleep, just hint me how I would get the area ;)

OpenStudy (samigupta8):

Heron's formula

OpenStudy (zyberg):

Thank you very much!

OpenStudy (samigupta8):

Welcum...:)